# Thompson sampling MAB multi-armed bandit

### v2

@@ Abstract
This paper focuses on selecting the arm with the highest variance from a set of $K$ independent arms. Specifically, we focus on two settings: (i) regret setting, that penalizes the number of pulls of suboptimal arms in terms of variance, and (ii) fixed-budget best arm identification (BAI) setting, that evaluates the ability of an algorithm to determine the arm with the highest variance after a fixed number of pulls. We develop a novel online algorithm called UCB-VV for the regret setting and show that its upper bound on regret for bounded rewards evolves as $\mathcal{O}\left(\log{n}\right)$ where $n$ is the horizon. By deriving the lower bound on the regret, we show that UCB-VV is order optimal. For the fixed budget BAI setting we propose the SHVV algorithm. We show that the upper bound of the error probability of SHVV evolves as $\exp\left(-\frac{n}{\log(K)H}\right)$, where $H$ represents the complexity of the problem, and this rate matches the corresponding lower bound. We extend the framework from bounded distributions to sub-Gaussian distributions using a novel concentration inequality on the sample variance. Leveraging the same, we derive a concentration inequality for the empirical Sharpe ratio (SR) for sub-Gaussian distributions, which was previously unknown in the literature. Empirical simulations show that UCB-VV consistently outperforms $\epsilon$-greedy across different sub-optimality gaps though it is surpassed by VTS, which exhibits the lowest regret, albeit lacking in theoretical guarantees. We also illustrate the superior performance of SHVV, for a fixed budget setting under 6 different setups against uniform sampling. Finally, we conduct a case study to empirically evaluate the performance of the UCB-VV and SHVV in call option trading on $100$ stocks generated using geometric Brownian motion (GBM).
Index Terms:
Multi-Armed bandits, risk, variance, regret, best arm identification.
@@ I Introduction and Motivation
Historically, variance has served as the first and foundational risk measure in portfolio optimization. The modern portfolio theory is based upon the classical Markowitz model, which also employs variance as a measure of risk [1, 2, 3]. In economic models, the decision-makers account for uncertainty and variance to maintain robustness [4]. In decision-making and optimization, identifying and mitigating sources of high variance is crucial to ensuring stability and robustness in outcomes [5, 6, 7]. Variance also plays a critical role in strategies involving trading call options, as higher variance increases the likelihood of the underlying asset’s price exceeding the strike price, thereby enhancing the intrinsic value of the option [8]. Also, it is a well-known fact that a call option on a volatile stock is more valuable than one that is less volatile [9].
Recent advances in risk-sensitive decision making involve more sophisticated risk measures such as value-at-risk (VaR) and conditional value-at-risk (CVaR) [10, 11, 12, 13, 14]. Unlike VaR and CVaR, which rely on tail-specific assumptions and involve complex computations related to probability thresholds and conditional expectations, variance is distribution-agnostic and computationally lightweight. From the perspective of selecting the arm with the highest variance, we study two settings of the multi-armed bandit (MAB) framework - regret minimization and BAI. Specifically, the regret setting attempts to minimize the number of pulls of sub-optimal arms in terms of the variance. On the contrary, the BAI setting attempts to determine the arm with the highest variance with a high probability.
For portfolio managers, maximizing the frequency of selecting the highest variance option unlocks opportunities to optimize returns while leveraging the asymmetrical payoff structure of call options from favorable price movements. For example, for a call option buyer, a higher variance is particularly advantageous as it increases the potential for substantial gains while capping the maximum loss to the premium paid for the option. In this paper, we investigate this aspect by developing bandit algorithms for call option trading using the variance associated with a stock. It must be noted that the results of our framework can be extended to the setting to pick the arm with the minimum variance as well.
Unlike the regret setting, the fixed-budget BAI focuses on identifying the best arm with minimum error for a fixed number of samples [15, 16, 17]. This is particularly relevant in scenarios like A/B testing, clinical trials, or recommendation systems, where the ultimate goal is to find the single best option. In finance, high-variance assets introduce instability and expose portfolios to erratic swings, requiring their identification and possible exclusion.
@@ I-A Related work
I-A1 Literature in regret and BAI setting
@@
The BAI problem has been extensively investigated across a broad range of settings, e.g, for linear bandits [22], cost-aware BAI [23], as well as risk-averse bandits [24]. In risk-averse BAI problem, [25] considers VaR-optimization. Also, [24], [26], optimizes a mean-CVaR framework within a fixed budget framework. They propose distribution oblivious algorithms, i.e, they do not require prior knowledge of the arms’ underlying distributions and are asymptotically near-optimal. However, these algorithms come with weaker performance guarantees compared to those that leverage prior knowledge of the arm distributions, highlighting an inherent trade-off between robustness and efficiency. Variance optimization, in particular, often necessitates stronger assumptions about the reward distributions (e.g., sub-Gaussianity) to ensure accurate and reliable estimation. Without such assumptions, variance estimates can become highly sensitive to noise and unreliable.
@@ I-A2 Literature in finance
In the context of finance, variance is used to model risk premium, providing insights into market sentiment and investor behavior. Specifically, the variance risk premium (a compensation that investors demand for bearing the risk associated with the volatility in the returns of an asset) plays a critical role in both theoretical and practical applications. This concept has been leveraged to develop trading strategies, such as selling volatility when the premium is high, as demonstrated in [27, 28]. Beyond its application in trading strategies, variance is also instrumental in understanding cross-market dynamics. For instance, Hamao et al. [29] investigated how volatility in one market, such as Japan, influences volatility in another, such as the U.S., using variance as a metric to assess volatility spillovers. Similarly, Engle et. al [30] examined intraday volatility in foreign exchange markets, employing variance to model the clustering of volatility over time. This clustering, a manifestation of the heteroscedastic behavior, is effectively captured by variance. Variance has been central to research on asset pricing and portfolio management. French et al. [31] explored the relationship between expected stock returns and volatility, with volatility quantified by variance, thereby informing optimal portfolio construction and asset pricing. Collectively, these studies illustrate the multifaceted role of variance in financial research and practice.
@@ I-B Contributions and Organization
    Regret algorithm and optimality: First, for the regret setting, we propose a UCB-type algorithm – UCB-VV that maximizes the frequency of selecting the arm with the highest variance. We show that the upper bound on the regret for UCB-VV is $\mathcal{O}(\log{n})$, and derive the corresponding information-theoretic lower bound to demonstrate that UCB-VV is order optimal. This is presented in Section II.
@@
    BAI algorithm and optimality: Next, we propose an arm-elimination based algorithm called SHVV to identify the arm with the highest variance within a fixed budget of pulls. Similar to the regret case, we derive the information-theoretic lower bound on the probability of error for the highest variance BAI and show that SHVV is order optimal. This is presented in Section III.
    Key novelty in 1) and 2): This is the first work that develops algorithms for selecting the arm with the highest variance and proves their order optimality.
@@
    Extension to sub-Gaussian distributions and applications to Sharpe ratio (SR): While the above results consider bounded distributions, we prove that our framework is more general and can be applied to sub-Gaussian distributions. For this, we derive a novel Bernstein-type concentration inequality for the sample variance of a sub-Gaussian distribution. Leveraging the same, we derive a novel concentration bound for the sample SR. We demonstrate the application of this result to devise a UCB-type algorithm to select the arm with the highest SR and provide a sketch for bounding its regret. This discussion is presented in Section IV.
@@
    Practical Application: Finally, following the theoretical development, in Section V, we compare the performance of UCB-VV with algorithms like $\epsilon$-greedy, VTS, and KL-UCB. We also demonstrate the superior performance of SHVV for a fixed budget setting under 6 different setups. Then, in Section VI, we validate the efficacy of the proposed algorithms (both UCB-VV and SHVV) on a real-world application – call option trading of GBM simulated $100$ stocks.
II Problem formulation for regret
Consider a MAB framework with $K\geq 2$ arms, a single player, and a time horizon $n$. We assume that the underlying rewards distributions of the arms are $\nu_{1},\nu_{2},\dots,\nu_{K}$, each of which is bounded [^1] In Section IV, we relax the boundedness assumption and demonstrate the extension of our framework to sub-Gaussian distributions. in $[l,u]$. The means, $\{\mu_{i}\}$, and the variances, $\{\sigma_{i}^{2}\}$, of the distributions are fixed but unknown to the player. When playing an arm $i$ at time $t$, a random reward $X_{i}(t)\sim\nu_{i}$ is received. The reward samples are independent across arms and time slots. A policy $\pi$ orchestrates the action of the player at each time instant and is driven by the historical data. Let $s_{i}(n)$ represent the number of plays of an arm $i$ until (and including) time $n$ by the player. In a standard stochastic bandit problem, the regret setting focuses on optimizing the trade-off between exploration and exploitation to maximize the expected cumulative reward over time. Without loss of generality, we assume that the optimal arm is unique and $\sigma^{2}_{\ast}=\max_{i\in[K]}\sigma_{i}^{2}$.
@@ II Problem formulation for regret
The biased empirical variance of an arm $i$ at time $t$, calculated from the i.i.d. samples $\{X_{i}(s_{i}(q))\}_{q=1}^{t}$ drawn from its distribution $\nu_{i}$ is given as
$\bar{V}_{i}(t)=\frac{1}{s_{i}(t)}\sum_{q=1}^{s_{i}(t)}\left(X_{i}(s_{i}(q))-\bar{\mu}_{i}(s_{i}(t))\right)^{2},$ (1)
where $\bar{\mu}_{i}(s_{i}(t))=\frac{1}{t}\sum_{q=1}^{t}X_{i}(s_{i}(q)).$
Then, for a policy $\pi$ and horizon $n$, we define the regret as
$\mathbb{E}[R(n)]=\sum_{i=1}^{K}\mathbb{E}[s_{i}(n)]\delta_{i},$ (2)
where $\delta_{i}=\sigma^{2}_{*}-\sigma_{i}^{2}$ denotes the sub-optimality gap of an arm $i$ with respect to the optimal arm.
Remark 1.
The regret defined in (2) deviates from the classical definition of regret (e.g., see [33]) that considers the regret as the expected deviation of the metric obtained by the algorithm from its optimal value. Unlike the mean, for maximizing the variance, the optimal policy may not be equivalent to sampling the arm with the highest variance at all times, but consists of strategic switching between the arms [34]. However, since our focus is the selection of the arm with the highest variance, and penalizing a policy each time it selects a variance sub-optimal arm, (2) is an appropriate choice of measuring the performance.
Note that employing a biased estimate of variance does not have any impact on our regret guarantees. As $t$ increases, the distinction between $t-1$ and $t$ becomes increasingly negligible due to their asymptotic equivalence. This adjustment impacts all arms equally without altering their relative ordering.
II-A UCB-VV Algorithm
@@
II-B Upper bound on regret of UCB-VV
First, we employ McDiarmid’s inequality to derive a concentration bound for the variance estimator $\bar{V}(n)$. Based on this, we derive an upper bound on regret of UCB-VV and show it evolves as $\mathcal{O}\log(n)$.
@@ II-B Upper bound on regret of UCB-VV
$\mathbb{P}\left(\left|\bar{V}(n)-\sigma^{2}\right|>\epsilon\right)\leq 2\exp\left(\frac{-2{n}\epsilon^{2}}{(u-l)^{2}}\right).$
@@
For a bandit setting with bounded rewards on $[l=0,u=1]$, the regret of UCB-VV is bounded as
$\text{R}_{\texttt{VV}}(n)\leq 8\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\frac{\log{n}}{\delta_{i}}+\left(1+\frac{\pi^{2}}{3}\right)\left(\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\delta_{i}\right).$
@@
See Appendix VII-A. ∎
Algorithm 1 UCB-VV
@@
II-C Lower bound on regret
@@
Definition 3.
A policy $\pi$ is $\alpha$-consistent $10<\alpha<$ 1, if for one parameter distributions and for all $i\neq*$, $\mathbb{E}[s_{i}(n)]\leq n^{\alpha},$ where * represents the optimal arm.
@@
We establish a preliminary lemma before proceeding to state the theorem.
@@
where constant $10<C_{2}<$ 1 is independent of $n$ and $\mathcal{F}$.
@@
See Appendix VII-B. ∎

### v3

@@ Abstract
This paper focuses on selecting the arm with the highest variance from a set of $K$ independent arms. Specifically, we focus on two settings: (i) misallocation minimization setting, that penalizes the number of pulls of suboptimal arms in terms of variance, and (ii) fixed-budget best arm identification setting, that evaluates the ability of an algorithm to determine the arm with the highest variance after a fixed number of pulls. We develop a novel online algorithm called UCB-VV for the misallocation minimization (MM) and show that its upper bound on misallocation for bounded rewards evolves as $\mathcal{O}\left(\log{n}\right)$ where $n$ is the horizon. By deriving the lower bound on the misallocation, we show that UCB-VV is order optimal. For the fixed budget best arm identification (BAI) setting we propose the SHVV algorithm. We show that the upper bound of the error probability of SHVV evolves as $\exp\left(-\frac{n}{\log(K)H}\right)$, where $H$ represents the complexity of the problem, and this rate matches the corresponding lower bound. We extend the framework from bounded distributions to sub-Gaussian distributions using a novel concentration inequality on the sample variance and standard deviation. Leveraging the same, we derive a concentration inequality for the empirical Sharpe ratio (SR) for sub-Gaussian distributions, which was previously unknown in the literature. Empirical simulations show that UCB-VV consistently outperforms $\epsilon$-greedy across different sub-optimality gaps though it is surpassed by VTS, which exhibits the lowest misallocation, albeit lacking in theoretical guarantees. We also illustrate the superior performance of SHVV, for a fixed budget setting under 6 different setups against uniform sampling. Finally, we conduct a case study to empirically evaluate the performance of the UCB-VV and SHVV in call option trading on $100$ stocks generated using geometric Brownian motion (GBM).
@@ I Introduction and Motivation
Historically, variance has served as the first and foundational risk measure in portfolio optimization. The modern portfolio theory is based upon the classical Markowitz model, which also employs variance as a measure of risk [21, 17, 24]. In economic models, the decision-makers account for uncertainty and variance to maintain robustness [14]. In decision-making and optimization, identifying and mitigating sources of high variance is crucial to ensuring stability and robustness in outcomes [34, 3, 25].
Recent advances in risk-sensitive decision making involve more sophisticated risk measures such as value-at-risk (VaR) and conditional value-at-risk (CVaR) [9, 8, 35, 36, 27]. Unlike VaR and CVaR, which rely on tail-specific assumptions and involve complex computations related to probability thresholds and conditional expectations, variance is distribution-agnostic and computationally lightweight. From the perspective of selecting the arm with the highest variance, we study two settings of the multi-armed bandit (MAB) framework - MM and BAI. Specifically, the MM setting attempts to minimize the number of pulls of sub-optimal arms in terms of the variance. On the contrary, the BAI setting attempts to determine the arm with the highest variance with a high probability. The fixed-budget BAI focuses on identifying the best arm with minimum error for a fixed number of samples [1, 6, 22]. This is particularly relevant in scenarios like A/B testing, clinical trials, or recommendation systems, where the ultimate goal is to find the single best option.
Variance has been central to research on asset pricing and portfolio management. High-variance assets introduce instability and expose portfolios to erratic swings, requiring their identification and possible exclusion. For portfolio managers, maximizing the frequency of selecting the highest variance option unlocks opportunities to optimize returns while leveraging the asymmetrical payoff structure of call options from favorable price movements. For example, for a call option buyer, a higher variance is particularly advantageous as it increases the potential for substantial gains while capping the maximum loss to the premium paid for the option. In this paper, we develop bandit algorithms for identification of arm with the highest variance in MM and BAI settings. It must be noted that the results of our framework can be extended to the setting to pick the arm with the minimum variance as well.
@@ I-A Related work
I-A1 Literature in bandit setting
@@
The BAI problem has been extensively investigated across a broad range of settings, e.g, for linear bandits [18], cost-aware BAI [31], as well as risk-averse bandits [20]. In risk-averse BAI problem, [10] considers VaR-optimization. Also, [20], [19], optimizes a mean-CVaR framework within a fixed budget framework. They propose distribution oblivious algorithms, i.e, they do not require prior knowledge of the arms’ underlying distributions and are asymptotically near-optimal. However, these algorithms come with weaker performance guarantees compared to those that leverage prior knowledge of the arm distributions, highlighting an inherent trade-off between robustness and efficiency. Variance optimization, in particular, often necessitates stronger assumptions about the reward distributions (e.g., sub-Gaussianity) to ensure accurate and reliable estimation. Without such assumptions, variance estimates can become highly sensitive to noise and hence unreliable.
@@ I-A2 Literature in finance
In the context of finance, variance is used to model risk premium, providing insights into market sentiment and investor behavior. Specifically, the variance risk premium (a compensation that investors demand for bearing the risk associated with the volatility in the returns of an asset) plays a critical role in both theoretical and practical applications. This concept has been leveraged to develop trading strategies, such as selling volatility when the premium is high, as demonstrated in [7, 38]. Beyond its application in trading strategies, variance is also instrumental in understanding cross-market dynamics. For instance, [13] investigated how volatility in one market, such as Japan, influences volatility in another, such as the U.S., using variance as a metric to assess volatility spillovers. Similarly, [11] examined intraday volatility in foreign exchange markets, employing variance to model the clustering of volatility over time. This clustering, a manifestation of the heteroscedastic behavior, is effectively captured by variance. [12] explored the relationship between expected stock returns and volatility, with volatility quantified by variance, thereby informing optimal portfolio construction and asset pricing. Variance also plays a critical role in strategies involving trading call options, as higher variance increases the likelihood of the underlying asset’s price exceeding the strike price, thereby enhancing the intrinsic value of the option [16]. Also, it is a well-known fact that a call option on a volatile stock is more valuable than one that is less volatile [29]. Collectively, these studies illustrate the multifaceted role of variance in financial research and practice.
@@ I-B Contributions and Organization
    Misallocation algorithm and optimality: First, for the MM setting, we propose an algorithm – UCB-VV that maximizes the frequency of selecting the arm with the highest variance. We derive an upper bound of order $\mathcal{O}(\log n)$ on the variance-misallocation of UCB-VV, together with a matching information-theoretic lower bound, thereby establishing the order optimality of the proposed algorithm for variance-based MM. These results are presented in Section II.
@@
    BAI algorithm and optimality: Next, we propose an arm-elimination based algorithm called SHVV to identify the arm with the highest variance within a fixed budget of pulls. Similar to the misallocation case, we derive matching upper and lower bounds on the probability of error for highest-variance identification and prove that SHVV is order optimal. These results are presented in Section III.
    Key novelty in 1) and 2): This is the first work that develops algorithms for selecting the arm with the highest variance and proves their order optimality. Derivation of all core probabilistic components for the variance objective is required even for BAI algorithms. Consequently, we resort to tools like McDiarmid’s inequality, Bernstein-type bound and KL-divergence for variance to establish sharp deviation guarantees for the sample variance.
@@
    Extension to sub-Gaussian distributions and applications to Sharpe ratio (SR): While the main theoretical results are derived for bounded rewards, we extend the framework to sub-Gaussian distributions. For this, we derive a novel Bernstein-type concentration inequality for the sample variance and standard deviation of a sub-Gaussian distribution. This extension is technically nontrivial because the square of a sub-Gaussian variable is sub-exponential, necessitating a two-term tail bound that cannot be obtained by direct extension of mean-based analyses. Leveraging this result, we derive a new, path-independent concentration inequality for the empirical SR statistic. We further demonstrate how this bound enables the construction of a UCB-type algorithm for SR maximization and provide a sketch of its regret analysis. These results are presented in Section IV.
@@
    Empirical evaluation and financial case study: Finally, following the theoretical development, in Section V, we empirically evaluate the proposed algorithms by comparing UCB-VV with baseline algorithms such as $\epsilon$-greedy and VTS. We also demonstrate the superior performance of SHVV for a fixed budget setting under 6 different setups. Then, in Section VI, we validate the efficacy of the proposed algorithms (both UCB-VV and SHVV) on a real-world application – we present a call option trading case study based on GBM-simulated stock prices as a qualitative financial illustration of variance driven decision making under convex payoffs.
II Problem formulation for misallocation
Consider a MAB framework with $K\geq 2$ arms, a single player, and a time horizon $n$. We assume that the underlying rewards distributions of the arms are $\nu_{1},\nu_{2},\dots,\nu_{K}$, each of which is bounded [^1] In Section IV, we relax the boundedness assumption and demonstrate the extension of our framework to sub-Gaussian distributions. in $[l,u]$. The means, $\{\mu_{i}\}$, and the variances, $\{\sigma_{i}^{2}\}$, of the distributions are fixed but unknown to the player. When playing an arm $i$ at time $t$, a random reward $X_{i}(t)\sim\nu_{i}$ is received. The reward samples are independent across arms and time slots. A policy $\pi$ orchestrates the action of the player at each time instant and is driven by the historical data. Let $s_{i}(n)$ represent the number of plays of an arm $i$ until (and including) time $n$ by the player. Without loss of generality, we assume that the optimal arm is unique and $\sigma^{2}_{\ast}=\max_{i\in[K]}\sigma_{i}^{2}$.
II-A Preliminaries for upper bound
@@
The unbiased empirical variance of an arm $i$ at time $t$, calculated from the i.i.d. samples $\{X_{i}(s_{i}(q))\}_{q=1}^{t}$ drawn from its distribution $\nu_{i}$ is given as
$\bar{V}_{i}(t)=\frac{1}{s_{i}(t)-1}\sum_{q=1}^{s_{i}(t)}\big(X_{i}(s_{i}(q))-\bar{\mu}_{i}(s_{i}(t))\big)^{2},$ (1)
where $\bar{\mu}_{i}(s_{i}(t))=\frac{1}{s_{i}(t)}\sum_{q=1}^{s_{i}(t)}X_{i}(s_{i}(q)).$
Definition 2.
Let $\sigma^{2}_{\ast}$ be the variance-optimal arm. The instantaneous loss at time $t$ is $\ell_{t}=\mathbb{I}\{I_{t}\neq*\}$. The cumulative misallocation up to horizon $n$ is
$\mathcal{M}(n)=\sum_{i\neq*}s_{i}(n)=\mathbb{E}\left[\sum_{t=1}^{n}\mathbb{I}\{I_{t}\neq*\}\right]$ (2)
The misallocation rate is
$\mathfrak{m}(n)=\frac{1}{n}\mathcal{M}(n)=\frac{1}{n}\,\mathbb{E}\left[\sum_{t=1}^{n}\mathbb{I}\{I_{t}\neq*\}\right]$
Unlike mean-reward maximization where the loss at time $t$ depends on the sub-optimality gap, variance maximization does not associate a meaningful incremental penalty with the magnitude of the variance gap. Once an arm is known to have lower variance than the optimal arm, any pull of that arm is equally undesirable, regardless of how small or large the gap may be. Consequently, the performance criterion naturally reduces to a $[0,1]$-type loss: at each time step, the learner either allocates the pull to the variance-optimal arm (no loss) or to a suboptimal arm (unit loss). This leads directly to the misallocation measure introduced in Definition 2, which evaluates how frequently the learner selects a suboptimal arm. This metric differs fundamentally from both cumulative regret which measures how bad each mistake is, and fixed-budget BAI which evaluates only the terminal recommendation. The misallocation rate instead captures the learner’s anytime behavior and penalizes every suboptimal selection during the learning process. It therefore provides an interpretable, gap-independent measure suited to the identification driven nature of variance maximization.
All algorithms and theoretical guarantees in this paper are based on the unbiased sample variance estimator defined in (1). Note that employing an biased estimate of variance does not have any impact on our guarantees. As $t$ increases, the distinction between t1t-1 and $t$ becomes increasingly negligible due to their asymptotic equivalence. This adjustment impacts all arms equally without altering their relative ordering.
II-B UCB-VV Algorithm
@@
II-C Upper bound on misallocation of UCB-VV
First, we employ McDiarmid’s inequality to derive a concentration bound for the variance estimator $\bar{V}(n)$. Based on this, we derive an upper bound on misallocation of UCB-VV and show it evolves as $\mathcal{O}\log(n)$.
@@ II-C Upper bound on misallocation of UCB-VV
$\mathbb{P}\left(\left|\bar{V}(n)-\sigma^{2}\right|>\epsilon\right)\leq 2\exp\left(\frac{-2{n}\epsilon^{2}}{(u-l)^{4}}\right).$
@@
For a bandit setting with bounded rewards on $[l=0,u=1]$, the misallocation of UCB-VV is bounded as
$\mathcal{M}_{\texttt{VV}}(n)\leq 8\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\frac{\log{n}}{\delta_{i}^{2}}+1+\frac{\pi^{2}}{3}.$
@@
See Appendix VIII-A. ∎
A key point to note is that while the algorithmic structure follows a UCB-like template leveraging confidence bounds to balance exploration and exploitation. However, the statistical analysis is specifically tailored to the variance objective.
Algorithm 1 UCB-$V$
@@
II-D Preliminaries for lower bound
@@
Definition 4.
A policy $\pi$ is $\alpha$-consistent $0<\alpha<1$, if for one parameter distributions and for all $i\neq*$, $\mathbb{E}[s_{i}(n)]\leq n^{\alpha},$ where * represents the optimal arm.
@@
where constant $0<C_{2}<1$ is independent of $n$ and $\mathcal{F}$.
@@
See Appendix VIII-B. ∎
II-E Lower bound on misallocation